The Gomez Company Inc bought a machine for $175,000. They paid 20% down and amortized the...

  1. The Gomez Company Inc bought a machine for $175,000. They paid 20% down and amortized the rest at 11.2% over a 3-year period.

    1. Find their Quarterly payment.

    2. Using MS-Excel, prepare an amortization table to show how the payments are made and how interest and principal balance are determined

How much will the company owe after 18 months? What is the amount of interest during their 5th payment?
Answer #1

Answer:

Loan amount = Machine cost - down payment = 175000 - 175000 * 20% = $140,000

Quarterly interest rate = 11.2% / 4 = 2.8%

Number of quarterly payments = 3 * 4 = 12

Quarterly payment = PMT (rate, nper, pv, fv, type) = PMT (2.8%, 12, -140000, 0, 0) = $13,897.30

Quarterly payment = $13,897.30

Amortization schedule is as follows:

A В C Loan Outstanding (Previous Quarter Ending D -C) $140,000.00 $130,022.70 Interest Amount Repayment (A- B) Quarterly (Pre

From above amortization schedule, we find:

Amount the company owe after 18 months (6 quarters) = $75,785.95

Amount of interest during their 5th payment = $2,754.73

The excel (amortization schedule) with 'show formula' is as below:

A. C В С Interest Amount (Previous Loan Outstanding (Previous Quarter Ending D- C) Quarterly Payment Repayment (A- B) End of